OpenStudy (anonymous):

Use the Binomial Theorem to find the coefficient of the x^8 y^3 term in the expansion of (2x^2+3y^3)^5 plzzzz show work thank you.

jimthompson5910 (jim_thompson5910):

n C r = (n!)/(r!(n-r)!) 5 C 1 = (5!)/(1!*(5-1)!) 5 C 1 = (5!)/(1!*4!) 5 C 1 = (5*4!)/(1!*4!) 5 C 1 = (5)/(1!) 5 C 1 = (5)/(1) 5 C 1 = (5)/(1) 5 C 1 = 5 ---------------------------------------------------------------------------------- (5 C 1)*(2^4)*(3^1) (5)*(2^4)*(3^1) (5)*(16)*(3) 240 So the coefficient of the x^8y^3 term is 240

OpenStudy (anonymous):

wow